A spring has a total length of 17.5 cm under a load of 250 g and 20.0cm under a load of 300 g. The extension of the spring per unit load is
1
Expert's answer
2020-07-02T17:14:39-0400

According to Hooke's law:


"F=k\Delta x=k(x_e-x_0),"

where "\Delta x" is elongation, "x_e" is the length of extended spring, "x_0" is the initial length. Hence, we can write two equations:


"m_1g=k(x_{e1}-x_0),\\
m_2g=k(x_{e2}-x_0).\\
0.25\cdot9.8=k(0.175-x_0),\\
0.3\cdot9.8=k(0.2-x_0).\\
k=19.6\text{ N/m},\\
x_0=5\text{ cm}."

The extension of the spring per unit load is, therefore


"\kappa=\frac{1}{k}=0.051\text{ m/N}."